OpenStudy (anonymous):

which of the following are the coordinates of the roots -4x^2-28x-48=0?

OpenStudy (anonymous):

\[-4x^2-28x-48=0?\]divide by -4 andget \[x^2+7x+12=0\] then maybe you can factor more easily to find the zeros

OpenStudy (anonymous):

I just need answer.

OpenStudy (anonymous):

-3 and -4

OpenStudy (anonymous):

use the discriminant to determine the nature of the roots of 4x^2+15x+10=0.

OpenStudy (anonymous):

Use Baskara: \[ax^2+bx+c=0\] \[x=\frac{ -b \pm \sqrt{b^2-4ac}}{ 2a }\]

OpenStudy (anonymous):

i do not understand that, at all..

OpenStudy (anonymous):

Its the formula used to solve second order equations, its usefull since its sinple, but for third and fourth order the formulas are too complicated, and for greater orders its impossible to find a formula. you get that by doing that: Multiply the equation by 4a \[4a^2x^2+4abx+4ac=0\]Sum b^2 \[4a^2x^2+4abx+4ac+b^2=0\]Wich is equal to \[(4ax+b)^2+4ac=0\]\[(2ax+b)^2=b^2-4ac\]\[2ax+b=\pm \sqrt{b^2-4ac}\]\[2ax=-b \pm \sqrt{b^2-4ac}\]\[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

Sorry for the mistake, after summing b^2 it should appear on both sides. But unintentionally I added it later on.